OpenStudy (anonymous):

? What is the solution of y − 4x = 0 and 3x + 6y = 9?

OpenStudy (anonymous):

A) x=0 , y=3/2 B) x=1/3 , y=4/3 C) x=1 , y=−1 D) x=1/4 , y=2/3

OpenStudy (pinkcandyrosez882):

Sorry idk! So sorry! Hope you get the answer soon!

OpenStudy (foodislife):

I think it's D. i'm not sure though